OpenStudy (anonymous):

Last ones! Simplify: -8(1-5x) (-7(5k-4) -5(7a-6) 5(n+6) 4(3r-8) 3(5+5x) (1+9x)x -10

OpenStudy (anonymous):

-8 + 40x -35k + 28 -35a + 30 5n +30 12r - 32 15 + 15x and the last one 9x^ + x - 10

OpenStudy (anonymous):

-8 + 40x -35k + 28 -35a + 30 5n +30 12r - 32 15 + 15x -10 -90x

OpenStudy (anonymous):

-8(1-5x) = -8 + 40x -7(5k-4) = -35k + 28 -5(7a-6) = -35a + 30 5(n+6) = 5n + 30 4(3r-8) = 12r - 32 3(5+5x) = 15 + 15x (1+9x)x -10 = x + 9x^2 - 10

OpenStudy (anonymous):

thank you soo much!